OpenStudy (anonymous):

Evaluate the expression when x = –2 and y = 3. 5x – 2y A. –18 B. 4 C. 24 D. –16

OpenStudy (***[isuru]***):

hi, the expression given is 5x - 2y and the given values r x = -2 y = 3 just plug those values in to that expression... [ 5x(-2) ] - [ 2x3] = -10 -6 = -16 hope this will help ya!!!